How many atoms of nitrogen are in 10.0g of N2O5 ?
1
Expert's answer
2020-05-03T15:12:24-0400

Solution:

Moles of N2O5 = Mass of N2O5 / Molar mass of N2O5

Molar mass of N2O5 is 108.0104 g/mol.

Then,

Moles of N2O5 = (10.0 g) / (108.0104 g/mol) = 0.09258 mol


N2O5 --> 2N

Moles of nitrogen (N) = 2 × Moles of N2O5 = (2 × 0.09258 mol) = 0.18516 mol


As 1 mole of any substance contains 6.022×1023 particles, 0.18516 mol of nitrogen contains:

Number of nitrogen atoms = (0.18516 mol) × (6.022×1023) = 1.115×1023 atoms = 1.1×1023 atomf of N.


Answer: 1.1×1023 atoms of nitrogen in 10.0g of N2O5.